What is the frequency of the note emitted by a disc siren with 30 holes, which is rotated at 500 revolutions per minute?!
1
Expert's answer
2012-09-04T10:25:16-0400
30 notes per revolution, 500 revolutions per minute:
N = 30 * 500 = 15000
notes per minute
N / 60 = 250 notes per second.
Answer: 250 Hz.